OpenStudy (anonymous):

Help real quick: 3x^3*2x^2

OpenStudy (anonymous):

its light stuff but I really forgot this -_-

OpenStudy (anonymous):

its been like 4 years I seen this lol..

OpenStudy (anonymous):

6x^5

OpenStudy (anonymous):

In this problem, you would multiply the coefficients, 3*2=6, and you would add the exponents x^3+x^2=x^5. So it would be 6x^5 :)